### v2

# Group Order is in $\mathrm{QCMA}$

Abstract
In this work, we show that verifying the order of a finite group given as a black-box is in the complexity class $\mathrm{QCMA}$. This solves an open problem asked by Watrous in 2000 in his seminal paper on quantum proofs and directly implies that the Group Non-Membership problem is also in the class $\mathrm{QCMA}$, which further proves a conjecture proposed by Aaronson and Kuperberg in 2006. Our techniques also give improved quantum upper bounds on the complexity of many other group-theoretical problems, such as group isomorphism in black-box groups.
1 Introduction
1.1 Background
QMA and Group Non-Membership.
The complexity class $\mathrm{QMA}$ (Quantum Merlin-Arthur) is one of the central complexity classes in quantum complexity theory. This class was first proposed by Knill [29] and Kitaev [28] as a natural quantum analogue of the classical class $\mathrm{NP}$ (or, more precisely, its randomized version called $\mathrm{MA}$), in which an all-powerful prover (named Merlin) sends a quantum proof to a verifier (named Arthur) who can perform bounded-error polynomial-time quantum computation. In 2000, Watrous [42] established its power by showing that several group-theoretic problems are in $\mathrm{QMA}$ in the black-box setting.
The concept of black-box group was first introduced (in the classical setting) by Babai and Szemerédi [12] to describe group-theoretic algorithms in the most general way, without depending on how elements are concretely represented and how group operations are implemented. In a black-box group, each group element is represented by a binary string and each group operation (group multiplication and inversion) is implemented using an oracle. Any efficient algorithm in the black-box group model thus gives rise to an efficient concrete algorithm when oracle operations can be replaced by efficient procedures, which can be done for many natural group representations, including permutation groups and matrix groups. In the quantum setting introduced by Watrous [42] and further investigated in several further works [2, 25, 30, 31, 43], the oracles should be able to handle quantum superpositions. Additionally, in the quantum setting, all these works assume that the group has unique encoding, i.e., each element should be encoded using a unique string (without unique encoding, even the most basic quantum primitives, such as computing the order of one element of the group, cannot be implemented).
The central problem considered in [42] is the Group Non-Membership problem defined below (where, for any elements $g_{1},\ldots,g_{k}$ of a group, we denote by $\langle g_{1},\ldots,g_{k}\rangle$ the subgroup generated by $g_{1},\ldots,g_{k}$):
Group Non-Membership
Instance: Group elements $g_{1},\ldots,g_{k}$ and $h$ in some finite group $\mathcal{G}$.
Question: Is $h$ outside the group generated by $g_{1},\ldots,g_{k}$ (i.e., is $h\notin G$ with $G=\langle g_{1},\ldots,g_{k}\rangle$) ?
Group Non-Membership is significantly more challenging than its complement, Group Membership, which asks if $h\in\langle g_{1},\ldots,g_{k}\rangle$: while Ref. [12] showed that Group Membership is in the class $\mathrm{NP}$, the best known classical upper bound for Group Non-Membership is $\mathrm{AM}$ (the class of problems that can be solved by a constant-round interactive proof system with public coins), by Babai [7]. We refer to Figure 1 for an illustration of the relations between the complexity classes discussed in this paper.
[…]
Figure 1: Known relations between the main complexity classes discussed in this paper. The inclusion $\mathrm{MA}\subseteq\Sigma_{2}^{\mathrm{P}}$ was shown by Babai [5]. The equality $\mathrm{QCMA}=\mathrm{QCMA}_{1}$ was shown by Jordan, Kobayashi, Nagaj and Nishimura [26]. All the other relations follow directly from the definitions.
Watrous [42] showed that Group Non-Membership is in $\mathrm{QMA}$. To prove this result, the quantum proof received from Merlin is the quantum superposition of all the elements in the group $G=\langle g_{1},\ldots,g_{k}\rangle$. Arthur checks that the quantum proof is valid (by checking that the quantum state is invariant under multiplication by $g_{1},\ldots,g_{k}$) and then checks that this quantum state is mapped to an orthogonal state when each element in the superposition is multiplied by $h$ (which guarantees that $h\notin G$). The key feature of this protocol is that it uses a quantum proof. Indeed, Watrous [42] also showed that there exist black-box groups for which Group Non-Membership is not in $\mathrm{MA}$, which shows that $\mathrm{QMA}$ is strictly more powerful than $\mathrm{MA}$ in the black-box setting (from a complexity-theoretic perspective, this can be interpreted as an oracle separation between $\mathrm{QMA}$ and $\mathrm{MA}$). Additionally, Watrous [42] showed that several additional group-theoretic problems (discussed later) are also contained in $\mathrm{QMA}$ via (fairly straightforward) reductions to Group Non-Membership. Besides Group Non-Membership being one of the most fundamental problems in $\mathrm{QMA}$, Watrous’ protocol is often used in educational material to illustrate the power of quantum proofs (see, e.g., [1, 38, 41, 46]), due to its simplicity. We also mention a later result by Grilo, Kerenidis and Sikora [20], which showed that Group Non-Membership is actually in $\mathrm{QMA}_{1}$, the one-sided version of $\mathrm{QMA}$.
Group Non-Membership and QCMA.
An important subclass of $\mathrm{QMA}$ is the class $\mathrm{QCMA}$ corresponding to problems where the proof is classical. One of the main open problems in quantum complexity theory, first posed by Aharonov and Naveh [3], is whether there exists a classical oracle separating $\mathrm{QMA}$ and $\mathrm{QCMA}$ (we refer to [2, 13, 19, 33, 34, 37, 47] for partial progress). In 2006, Aaronson and Kuperberg [2] showed that Group Non-Membership is actually in the class $\mathrm{QCMA}$ under some group-theoretic assumptions, which gives evidence that Group Non-Membership is not a good candidate for a separation between $\mathrm{QMA}$ and $\mathrm{QCMA}$. Aaronson and Kuperberg further conjectured that Group Non-Membership is actually in $\mathrm{QCMA}$ unconditionally:
Conjecture 1 ([2]).
Group Non-Membership is in $\mathrm{QCMA}$.
No progress has been made on this conjecture since 2006.
Group Order Verification.
As explained above, Group Non-Membership is a fundamental task in group theory. An even more powerful primitive is computing the order of a group. For a group $G$, we write its order (i.e., the number of elements in $G$) as $|G|$. We introduce the decision version of this problem as follows:
Group Order Verification
Instance: Group elements $g_{1},\ldots,g_{k}$ in some finite group $\mathcal{G}$, and a positive integer $m$.
Question: Is the order of the group generated by $g_{1},\ldots,g_{k}$ equal to $m$ (i.e., is $|G|=m$
                                 with $G=\langle g_{1},\ldots,g_{k}\rangle$) ?
Group Non-Membership reduces to Group Order Verification since $h\notin G$ if and only if $|G|\neq|\langle g_{1},\ldots,g_{k},h\rangle|$. [^1] Note that this reduction is nondeterministic: it assumes the existence of a prover who can “guess” the orders of the two groups $G$ and $\langle g_{1},\ldots,g_{k},h\rangle$, which can then be verified using a protocol for Group Order Verification. Such a nondeterministic reduction will be enough since in this paper we only consider complexity classes with a prover. The best known upper bound on Group Order Verification is $\mathrm{AM}\cap\mathrm{coAM}$, by Babai [7]. Since Group Non-Membership, which belongs to $\mathrm{QMA}$, reduces to Group Order Verification, this leads to one of the main open problems proposed in [42]:
Open Problem 1 ([42]).
Is Group Order Verification in $\mathrm{QMA}$?
No progress has been made on this problem since 2000.
1.2 Our results
Statement of our results.
In this paper, we prove 1 and solve 1. Here is our main result:
Theorem 1.1.
Group Order Verification is in $\mathrm{QCMA}$.
Theorem 1.1 solves 1. Our result is actually significantly stronger: it shows that Group Order Verification is not only in $\mathrm{QMA}$, but also in $\mathrm{QCMA}$. As observed in [7], an upper bound on the complexity of Group Order Verification leads to the same upper bound for the complement: in order to verify that $|G|\neq m$, Merlin can send the true order of $G$ and then Arthur can use the protocol of Theorem 1.1 for checking whether it is really the true order and differs from $m$. We thus obtain the following stronger statement:
Corollary 1.1.
Group Order Verification is in $\mathrm{QCMA}\cap\mathrm{coQCMA}$.
Since Group Non-Membership reduces to Group Order Verification, as another immediate corollary, we obtain a proof of 1:
Corollary 1.2.
Group Non-Membership is in $\mathrm{QCMA}$.
Table 1: This table compares our new upper bounds with the upper bounds from the literature.
Problem Prior upper bounds New quantum
Classical Quantum upper bound
Group Order Verification $\mathrm{AM}\cap\mathrm{coAM}$    [7] - $\mathrm{QCMA}\cap\mathrm{coQCMA}$ (Cor. 1.1)
Group Non-Membership $\mathrm{AM}\cap\mathrm{coNP}$ [7, 12] $\mathrm{QMA}$     [42] $\mathrm{QCMA}$         (Cor. 1.2)
Group Isomorphism $\mathrm{AM}\cap\Sigma_{2}^{\mathrm{P}}$     [7, 12] - $\mathrm{QCMA}$         (Cor. 1.3)
Homomorphism, $\mathrm{AM}\cap\mathrm{coAM}$    [7] - $\mathrm{QCMA}\cap\mathrm{coQCMA}$ (Cor. 1.4)
Minimal Normal Subg.
Proper Subgroup $\mathrm{QMA}$     [42]
Simple Group $\mathrm{coQMA}$[42]
Intersection, Centralizer, $\mathrm{coQCMA}$       (Cor. 1.5)
Maximal Normal Subg.
Other than Group Order Verification and Group Non-Membership, we obtain new quantum upper bounds for the complexity of many group-theoretic problems: Group Isomorphism, Homomorphism, Minimal Normal Subgroup, Proper Subgroup, Simple Group, Intersection, Centralizer and Maximal Normal Subgroup (the formal definition of these problems is given in Section 6). These eight problems have been considered in the classical setting in [7, 12]. The last five problems have been considered in the quantum setting in [42]. By combining Corollary 1.1 with the proof techniques from [7, 42], we easily obtain the following results:
Corollary 1.3.
Group Isomorphism is in $\mathrm{QCMA}$.
Corollary 1.4.
Homomorphism, Minimal Normal Subgroup, Proper Subgroup and Simple Group are in $\mathrm{QCMA}\cap\mathrm{coQCMA}$.
Corollary 1.5.
Intersection, Centralizer and Maximal Normal Subgroup are in $\mathrm{coQCMA}$.
All the results are summarized in Table 1.
Related work.
When writing this paper, we learned from Michael Levet [32] and James Wilson [44] that Alexander Hulpke, Martin Kassabov, Ákos Seress and James Wilson have obtained a proof of the existence of a short presentation for the Ree groups of rank one. The proof, which is 60-page long, is unpublished (and not expected to be published). The existence of such a short presentation leads to an alternative way of proving Theorem 4.1, by using Proposition 2.1 instead of our isomorphism test.
1.3 Overview of the proof strategy
We give below an overview of the strategy we use to prove Theorem 1.1.
Let us first describe some basic notation and notions of group theory — more details are given in Section 2.2. For a group $G$, we write $H\leq G$ (resp., $H\unlhd G$) to express that $H$ is a subgroup (resp., normal subgroup) of $G$. We denote by $\{e\}$ the trivial subgroup of $G$. A composition series of a group $G$ is a decomposition of the group into simple groups (a simple group is a nontrivial group that has no nontrivial normal subgroup and thus cannot be further decomposed), which are called the composition factors of $G$. The “classification theorem of finite simple groups” states that every finite simple group belongs to one of 18 infinite families of simple groups, or is one of 26 sporadic simple groups. As a consequence, each simple group can be described by a short string called its standard name.
Babai-Beals filtration.
The starting point of our strategy is the Babai-Beals filtration. Babai and Beals [8] showed that any group $G$ has a decomposition
$\{e\}\unlhd\mathrm{Sol}(G)\unlhd\mathrm{Soc^{*}}(G)\unlhd\mathrm{Pker}(G)\unlhd G\,$,
where $\mathrm{Sol}(G)$ and $\mathrm{Pker}(G)$ are two normal subgroups of $G$ called the solvable radical and the permutation kernel, respectively, and $\mathrm{Soc^{*}}(G)$ is another normal subgroup (all these subgroups are defined in Section 3.1, but their definition is not needed for this overview). Ref. [8] showed that in randomized polynomial time, it is possible to compute a set of generators for $\mathrm{Pker}(G)$. Additionally, given a set of generators for $\mathrm{Pker}(G)$, it is possible in deterministic polynomial time to test membership in $\mathrm{Pker}(G)$ and compute the order $|G/\mathrm{Pker}(G)|$. Since
$|G|=|\mathrm{Pker}(G)|\cdot|G/\mathrm{Pker}(G)|,$
in order to compute $|G|$ we thus only need to compute the order of $\mathrm{Pker}(G)$.
While how to compute efficiently $\mathrm{Sol}(G)$ and $\mathrm{Soc^{*}}(G)$ is unknown (even with the help of a prover), these two subgroups have an important property: $\mathrm{Sol}(G)$ and $\mathrm{Pker}(G)/\mathrm{Soc^{*}}(G)$ are solvable (a solvable group is a group that is “not too much non-abelian,” in the sense that all its composition factors are cyclic). Note that while the solvability of $\mathrm{Sol}(G)$ is easy to show, the solvability of $\mathrm{Pker}(G)/\mathrm{Soc^{*}}(G)$ is based on Schreier conjecture, which was proposed by Schreier in 1926, and is now known to be true as a result of the classification of finite simple groups (no simpler proof is known).
Stategy to compute $|\mathrm{Pker}(G)|$.
In order to compute the order of $\mathrm{Pker}(G)$, we first observe that the Babai-Beals filtration implies the existence of a solvable subgroup $H_{0}$ and $2s$ elements $\beta_{1},\ldots,\beta_{s},\gamma_{1},\ldots,\gamma_{s}\in\mathrm{Pker}(G)$ such that, when defining
$H_{i}$ $=\langle H_{0},\beta_{1},\ldots,\beta_{i},\gamma_{1},\ldots,\gamma_{i}\rangle$
for each $i\in[s]$, the chain of inclusions
$\{e\}\unlhd H_{0}\unlhd H_{1}\unlhd\cdots\unlhd H_{s}\unlhd\mathrm{Pker}(G)$ (1)
holds, where $\mathrm{Pker}(G)/H_{s}$ is solvable and $H_{i}/H_{i-1}$ is a simple group for each $i\in[s]$. [^2] For instance, we can take $H_{0}=\mathrm{Sol}(G)$ and $H_{s}=\mathrm{Soc^{*}}(G)$ to show the existence of such a decomposition. We nevertheless do not require the conditions $H_{0}=\mathrm{Sol}(G)$ and $H_{s}=\mathrm{Soc^{*}}(G)$ since they cannot be checked efficiently (as mentioned above, we do not know how to efficiently compute $\mathrm{Sol}(G)$ and $\mathrm{Soc^{*}}(G)$). Observe that
$|\mathrm{Pker}(G)|=|H_{0}|\cdot|H_{1}/H_{0}|\cdot|H_{2}/H_{1}|\cdots|H_{s}/H_{s-1}|\cdot|\mathrm{Pker}(G)/{H_{s}}|$
holds. Here $|H_{0}|$ and $|\mathrm{Pker}(G)/H_{s}|$ are easy to compute with the help of Merlin since they are solvable. [^3] Actually, $|H_{0}|$ can be computed even without Merlin’s help by using Watrous’ algorithm for solvable groups [43]. It thus remains to compute $|H_{i}/H_{i-1}|$ for each $i\in[s]$.
While it is unknown whether the decomposition (1) can be computed in polynomial time, we can ask Merlin to “guess” it and send it to Arthur. Concretely, Merlin sends a set of generators for each subgroup $H_{i}$ and the standard name of the simple group $H_{i}/H_{i-1}$. The standard names enable Arthur to learn each $|H_{i}/H_{i-1}|$, and thus to compute $|\mathrm{Pker}(G)|$. A dishonest Merlin, however, might cheat and send a wrong standard name or can even send a series in which some $H_{i}/H_{i-1}$ is not simple. The main obstacle is thus to check that $H_{i}/H_{i-1}$ is really isomorphic to the simple group specified by Merlin (by its standard name).
Isomorphism test.
One promising strategy for testing if a group $\Sigma$ is isomorphic to a known (not necessarily simple) group $S$ is to use a randomized homomorphism test. A similar strategy was also used by Aaronson and Kuperberg to analyze the query complexity of Group Non-Membership [2].
Let $s_{1},\ldots,s_{k}$ be a set of generators of $S$ known to both Arthur and Merlin. We ask Merlin to send elements $g_{1},\ldots,g_{k}\in\Sigma$. If $\Sigma\cong S$ and Merlin is honest, he will send $g_{i}=\phi(s_{i})$ for each $i\in[k]$, for some isomorphism $\phi\colon S\to\Sigma$. For the checking procedure, Arthur defines a map $f\colon S\to\Sigma$ by extending the partial map $s_{i}\mapsto g_{i}$ into a map on all $S$ as if it were a homomorphism. For instance, for an element $s\in S$ that can be written as $s=s_{1}s_{2}s_{1}s_{3}$, Arthur will set $f(s)=g_{1}g_{2}g_{1}g_{3}$. Arthur then takes two elements $s$ and $s^{\prime}$ uniformly at random in $S$ and checks if
$f(ss^{\prime})=f(s)f(s^{\prime})$ (2)
holds. By standard results on property testing (e.g., [14]), we can show that passing this test with high probability guarantees that there exists a homomorphism from $S$ to $\Sigma$.
To be successful, this approach has to satisfy three important requirements:
  • A.
    Arthur needs to be able to efficiently represent an arbitrary element $s\in S$ as a product of elements from the fixed set $\{s_{1},\ldots,s_{k}\}$. This representation should also be unique for $f$ to be well-defined.
  • B.
    Arthur needs to be able to efficiently check that the homomorphism whose existence is guaranteed when passing the homomorphism test is actually an isomorphism, i.e., a bijection.
  • C.
    Arthur needs to be able to efficiently check if (2) holds.
The first two requirements were also mentioned by Aaronson and Kuperberg [2] as obstacles to prove that Group Non-Membership is in $\mathrm{QCMA}$. In particular, Task B was handled in [2] by solving an instance of the Normal Hidden Subgroup Problem using the quantum algorithm by Ettinger, Høyer and Knill [18], which has polynomial query complexity but in general exponential time complexity.
Note that Aaronson and Kuperberg [2] applied the homomorphism test to the whole group $\Sigma=G$. In our strategy, however, we are working on a composition factor $\Sigma=H_{i}/H_{i-1}$, i.e., we only need to consider the case where $S$ is a simple group. For simple groups, Tasks A and B can be implemented efficiently. For Task A, simple groups have a concrete representation for which we can efficiently represent any element as a product of elements from the fixed set (this is nontrivial and requires advanced techniques, such as the machinery for matrix groups developed by Babai, Beals and Seress [9]). For Task B, we can fairly easily guarantee that the homomorphism is a bijection by exploiting the property that simple groups do not have nontrivial subgroups (another interpretation is that the Normal Hidden Subgroup Problem is easy in simple groups since the only normal subgroups are the trivial subgroup and the whole group). In other words, we are able to bypass the first two obstacles because we are working on composition factors, and not on the whole group.
The price to pay is that Task C now becomes very challenging. When working on the whole group $\Sigma=G$ as done in [2], checking if (2) holds is trivial since the oracle for the black-box group $\mathcal{G}$ can be directly applied. When $\Sigma=H_{i}/H_{i-1}$, this is not the case anymore: checking if (2) holds is equivalent to checking if
$f(ss^{\prime})f(s^{\prime})^{-1}f(s)^{-1}\in H_{i-1}$
holds, which requires the ability to check membership in $H_{i-1}$. This is challenging since the group $H_{i-1}$ can be arbitrary. Note that it is crucial that the elements $s$ and $s^{\prime}$ chosen by Arthur are hidden from Merlin, otherwise Merlin can cheat by choosing $g_{i}$’s such that Eq. 2 holds (only) for those specific $s$ and $s^{\prime}$. For this reason, we cannot use Merlin to directly help Arthur check membership in $H_{i-1}$ (e.g., by sending a membership certificate). Instead, Arthur should be able to efficiently test membership in $H_{i-1}$ by himself. This is the main difficulty we have to overcome in this work.
Replacing membership in $H_{i-1}$ by membership in $H_{0}$.
We show (in Theorem 3.1) the following crucial consequence of the Babai-Beals filtration: there exists a decomposition of the form (1) that satisfies the additional condition
$H_{i}/H_{i-1}\cong\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}\,\textrm{, for all }i\in[s].$
In our protocol for Group Order Verification (described in Section 5), an honest Merlin sends a decomposition satisfying $(\star)$. In order to check if $H_{i}/H_{i-1}\cong S_{i}$ for some specific simple group $S_{i}$, it is thus enough to check if $\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}\cong S_{i}$. As explained above, to use the homomorphism test, we need to be able to efficiently check membership in $H_{0}$. Crucially, $H_{0}$ is now a solvable group, and we can thus use Watrous’ polynomial-time quantum algorithm for membership testing in solvable groups [43] to implement the homomorphism test efficiently.
In the case of a dishonest Merlin, we can still guarantee that the decomposition (1) satisfies $\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}\cong S_{i}$ for all $i\in[s]$, [^4] A decomposition such that $\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}\cong S_{i}$ for all $i\in[s]$ is called a nice decomposition in Section 3 (see Definition 3.1). but cannot guarantee that it satisfies Condition ($\star$). We are nevertheless able to show (in Proposition 3.1) that $\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}\cong S_{i}$ implies that $|H_{i}/H_{i-1}|$ is a divisor of $|S_{i}|$, and show that guaranteeing that $|H_{i}/H_{i-1}|$ is a divisor of $|S_{i}|$ is enough for our purpose.
In order to finish establishing the soundness of the protocol, further work is needed. We should especially deal with the potential cheating strategy in which Merlin sends, instead of (1), the chain of subgroup
$\{e\}\unlhd H_{0}\unlhd H_{1}\unlhd\cdots\unlhd H_{s}\unlhd K$ (2)
for a proper subgroup $K\lneq\mathrm{Pker}(G)$. To prevent such a cheating, we observe that $K=\mathrm{Pker}(G)$ if and only if the composition factors of $G/\mathrm{Pker}(G)$ match (with multiplicity) the composition factors of $G/K$. We then use another deep property of the Babai-Beals filtration: $G/\mathrm{Pker}(G)$ is isomorphic to a symmetric group of small degree, which implies that the composition factors of $G/\mathrm{Pker}(G)$ are fairly “easy”. We can thus check if the composition factors of $G/\mathrm{Pker}(G)$ match (with multiplicity) the composition factors of $G/K$ fairly easily with the help of Merlin.
The Ree groups of rank one.
Instead of testing if $\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}\cong S_{i}$ for each $i\in[s]$ using the isomorphism test described above, we observe that we actually only need to do it for one class of simple groups: the Ree groups of rank one. For all the other simple groups, we can use a simpler approach, already proposed in [8], based on the existence of short presentations (see Proposition 2.1). In this paper, we thus describe the isomorphism test only for the Ree groups of rank one (in Section 4).
2 Preliminaries
In this section, we describe the notions of complexity theory, group theory and black-box groups needed to show our results. For any positive integer $s$, we write $s]=\{1,\ldots,s\}$.
2.1 Quantum complexity theory
We assume that the reader is familiar with the most basic concepts and terminology of quantum computing, such as quantum circuits and measurements. The main technical contribution of this work is to construct classical certificates for order verification that can be checked by known quantum algorithms (e.g., Watrous’ quantum algorithms for solvable groups [43]). The claims of this paper can be verified without further expertise in quantum computing if the reader is willing to consider these quantum algorithms as black-boxes. We just give below the formal definition of the complexity class $\mathrm{QCMA}$.
Definition 2.1.
A problem $A=(A_{{\rm yes}},A_{{\rm no}})$ is in $\mathrm{QCMA}$ if there is a polynomial-time quantum algorithm $V$ (by a verifier called Arthur) such that:
Completeness
For any $x\in A_{{\rm yes}}$, there is a polynomial-length binary string $w_{x}$ called a certificate (from a prover called Merlin) such that $V$ accepts on input $(x,w_{x})$ with probability at least $2/3$.
Soundness
For any $x\in A_{{\rm no}}$, $V$ accepts with probability at most $1/3$ on input $(x,w)$ for any polynomial-length binary string $w$.
The quantum algorithm $V$ with certificates $\{w_{x}\}_{x\in A_{{\rm yes}}}$ is called a $\mathrm{QCMA}$ protocol.
2.2 Group theory
A group $G$ is called solvable if there exist $g_{1},\cdots,g_{s}\in G$ such that when defining $H_{i}=\langle g_{1},\ldots,g_{i}\rangle$ for each $i\in[s]$,
$\{e\}=H_{0}\unlhd H_{1}\unlhd\cdots\unlhd H_{s}=G.$
Note that $H_{i}/H_{i-1}$ is necessarily cyclic in this case, for each $i\in[s]$. For any integer $k\geq 1$, we denote by $\mathrm{Sym}(k)$ the symmetric group of degree $k$. Let $S$ be a set of generators of the group $G$. We call a sequence $(g_{1},\ldots,g_{t})$ of elements of $G$ a straight-line program over $S$ if each $g_{i}$ is either a member of $S$ or an element of the form $g_{j}^{-1}$ or $g_{j}g_{k}$ from some $j,k<i$. The length of the straight-line program is $t$. The element reached by the straight-line program is the last element $g_{t}$.
We will use the following easy fact in Section 5.
Fact 1.
(see, e.g., [39, Proposition 5.42]) For any finite group $G$ and any prime power $p^{t}$, $p^{t}$ divides $|G|$ if and only if $G$ has a subgroup of order $p^{t}$.
Composition series and simple groups.
A simple group is a nontrivial group with no nontrivial normal subgroup. Any group can be decomposed into simple groups via composition series.
Definition 2.2.
Let $G$ be a finite group. A composition series of $G$ is a list of subgroups $H_{0},H_{1},\ldots,H_{s}$, for some integer $s$, such that
  • (a)
    $\{e\}=H_{0}\unlhd H_{1}\unlhd\cdots\unlhd H_{s}=G$ ;
  • (b)
    the quotient group $H_{i}/H_{i-1}$ is a simple group for each $i\in[s]$.
The composition factors of $G$ are the quotients $H_{1}/H_{0},H_{2}/H_{1},\ldots,H_{s}/H_{s-1}$.
Any finite group has a composition series (with $s=O(\log|G|)$). While a group may have more than one composition series, the Jordan-Hölder theorem (e.g., [17, Theorem 22]) shows that they have the same length and the same composition factors, up to permutation and isomorphism.
The “classification theorem of finite simple groups” (see, e.g., [45]) is a theorem which states that every finite simple group belongs to one of 18 infinite families of simple groups (each family being indexed by one or two parameters), or is one of 26 sporadic simple groups. This gives $18+26=44$ types of finite simple groups. Each finite simple group $S$ can thus be represented by a binary string $z=(z_{1},z_{2})$ where $z_{1}$ is a constant-length binary string indicating its type and $z_{2}$ is a $O(\log|S|)$-length string representing its parameters ($z_{2}$ is empty if $S$ is a sporadic simple group). We call $z$ the standard name of the finite simple group $S$. Conversely, given a binary string $z$ corresponding to a standard name of a finite simple group, we write $\mathrm{gr}(z)$ the simple group represented by $z$. Given a standard name $z$, the order $|\mathrm{gr}(z)|$ can be easily computed.
The Ree groups of rank one.
The family of Ree groups of rank one is among the least understood families of finite simple groups. In this paper, we define these groups by using their natural matrix representation, and use the same set of generators as in [4, 27].
The family of Ree groups of rank one is indexed by a positive integer $a$. Write $q=3^{2a+1}$ and $t=3^{a}$. Let $\mathbb{F}_{q}$ be the finite field of order $q$ and $\omega$ be a primitive element of $\mathbb{F}_{q}$. Consider the group $\mathrm{GL}(7,q)$ of invertible matrices of dimension 7 over $\mathbb{F}_{q}$. The Ree group of rank one, which we denote by $\mathrm{R}(q)$, [^5] This group is also written as $^{2}\mathrm{G}_{2}(q)$ in the literature. is the subgroup of $\mathrm{GL}(7,q)$ generated by the following three matrices:
$\Gamma_{1}$ $=\begin{bmatrix}1&1&0&0&-1&-1&1\\ 0&1&1&1&-1&0&-1\\ 0&0&1&1&-1&0&1\\ 0&0&0&1&1&0&0\\ 0&0&0&0&1&-1&1\\ 0&0&0&0&0&1&-1\\ 0&0&0&0&0&0&1\end{bmatrix},\hskip 71.13188pt\Gamma_{2}=\begin{bmatrix}0&0&0&0&0&0&-1\\ 0&0&0&0&0&-1&0\\ 0&0&0&0&-1&0&0\\ 0&0&0&-1&0&0&0\\ 0&0&-1&0&0&0&0\\ 0&-1&0&0&0&0&0\\ -1&0&0&0&0&0&0\end{bmatrix}$,
$\Gamma_{3}$ $=\begin{bmatrix}\omega^{t}&0&0&0&0&0&0\\ 0&\omega^{1-t}&0&0&0&0&0\\ 0&0&\omega^{2t-1}&0&0&0&0\\ 0&0&0&1&0&0&0\\ 0&0&0&0&\omega^{1-2t}&0&0\\ 0&0&0&0&0&\omega^{t-1}&0\\ 0&0&0&0&0&0&\omega^{-t}\end{bmatrix}$.
The order of $\mathrm{R}(q)$ is $q^{3}(q^{3}+1)(q-1)$.
Short presentations of groups.
A presentation of a group $G$ is a definition of $G$ in terms of generators and relations (see, e.g., [17]). The length of the presentation is the total number of characters required to write down all relations between the generators. It is known that most simple groups have short presentations:
Theorem 2.1 ([10, 24]).
All the finite simple groups, with the possible exception of the family of Ree groups of rank one, have a polylogarithmic-length presentation (i.e., a presentation of length polynomial in the logarithm of the order of the group). Moreover, these polylogarithmic-length presentations can be efficiently computed from the standard name of the simple group.
In particular, any composition factor of a solvable group has a polylogarithmic-length presentation. For any group isomorphic to a subgroup of a permutation group of small degree, any composition factor also has a polylogarithmic-length presentation. This result is well-established in the literature (see, e.g., [35, Section 6]), but not explicitly stated. We provide a proof here for completeness.
Theorem 2.2.
If $G\leq\mathrm{Sym}(k)$, then any composition factor of $G$ has a $\mathrm{poly}(k)$-length presentation.
Proof.
Let $\{e\}=H_{0}\unlhd H_{1}\unlhd\cdots\unlhd H_{s}\leq G$ be a composition series of $G$. Each $H_{i-1}$ is a maximal normal subgroup in $H_{i}$ and $H_{i}/H_{i-1}$ is a simple group. Then for each $i\in[s]$, $H_{i}/H_{i-1}$ is isomorphic to a subgroup of $\mathrm{Sym}(k)$ (see, e.g, [23]).
By Theorem 2.1, $H_{i}/H_{i-1}$ has a presentation of length $\mathrm{poly}(k)$ except when $H_{i}/H_{i-1}\cong\mathrm{R}(q)$ is a Ree group of rank one.
Suppose $H_{i}/H_{i-1}\cong\mathrm{R}(q)$. It is known that $(q^{3}+1)$ is the least positive integer such that there is an injective homomorphism $\phi:\mathrm{R}(q)\to\mathrm{Sym}(q^{3}+1)$ (see, e.g., [21, Table 4]). Therefore, $|H_{i}/H_{i-1}|=|\mathrm{R}(q)|=q^{3}(q^{3}+1)(q-1)\leq(q^{3}+1)^{3}\leq k^{3}$. Taking the generating set as the set of all elements of $H_{i}/H_{i-1}$ and the set of relations as all the multiplication relations in the multiplication table, we get a $\mathrm{poly}(k)$-length presentation for $H_{i}/H_{i-1}\cong\mathrm{R}(q)$. ∎
Homomorphism test.
Testing the homomorphism of finite groups is a well-studied topic in property testing. In this paper we will use the following result from [2], which is based on [14].
Lemma 2.1 (Propositions 5.2 and 5.3 in [2]).
Let $G,G^{\prime}$ be two groups and consider a function $f\colon G\to G^{\prime}$. Assume that the inequality
$\Pr_{r_{1},r_{2}\in G}\>[f(r_{1}r_{2})=f(r_{1})f(r_{2})]\geq 9/10$ (3)
holds. Then there exists a unique homomorphism $\phi\colon G\to G^{\prime}$ such that
$\Pr_{x\in G}\>[f(x)\neq\phi(x)]\leq 1/10.$
2.3 Black-box groups
A black-box group is a representation of a group $\mathcal{G}$ introduced by Babai and Szemerédi [12] in which each element of $\mathcal{G}$ is encoded by a binary string of a fixed length $n=O(\log|\mathcal{G}|)$. Let $s\colon\mathcal{G}\to\{0,1\}^{n}$ denote the encoding of elements as binary strings. If $s$ is injective, we say that $\mathcal{G}$ is a black-box group with unique encoding. If the encoding is not unique, an oracle for identity testing (i.e., deciding whether or not a given string encodes the identity element of $\mathcal{G}$) is available.
In the classical setting, classical oracles are available to perform group operations (each call to the oracles can be done at unit cost). A first oracle performs the group product: given two strings representing two group elements $g$ and $h$, the oracle outputs the string representing $gh$. A second oracle performs inversion: given a string representing an element $g\in\mathcal{G}$, the oracle outputs the string representing the element $g^{-1}$. These two oracles output an error message on non-valid inputs (i.e., strings in $\{0,1\}^{n}\setminus s(\mathcal{G})$). All the classical algorithms and protocols discussed in this paper do not require that $\mathcal{G}$ has unique encoding.
In the quantum setting, the oracles performing the group operations have to be able to deal with quantum superpositions. As in prior works [2, 25, 30, 31, 42, 43], in the quantum setting we always consider black-box groups with unique encoding. Let $s\colon\mathcal{G}\to\{0,1\}^{n}$ denote the injective encoding of elements as binary strings. Two quantum oracles are available (each call to the oracles can be done at unit cost). The first oracle maps $|s(g)\rangle|s(h)\rangle$ to $|s(g)\rangle|s(gh)\rangle$, for any two elements $g,h\in\mathcal{G}$. The second quantum oracle maps $|s(g)\rangle|s(h)\rangle$ to $|s(g)\rangle|s(g^{-1}h)\rangle$, for any $g,h\in\mathcal{G}$. These two oracles output an error message on non-valid inputs (in the quantum setting this is implemented by introducing a third 1-qubit register that is flipped when the inputs are not valid — see [42] for details).
We describe below several classical and quantum techniques for black-box groups.
Testing solvability and approximate sampling in black-box groups.
We will use the following classical randomized algorithms from [6, 8] to sample nearly uniformly elements in black-box groups and test solvability.
Theorem 2.3 ([6]).
Let $\mathcal{G}$ be a black-box group. For any $G\leq\mathcal{G}$ and any $\varepsilon>0$, there exists a classical randomized algorithm running in time polynomial in $\log(|\mathcal{G}|)$ and $\log(1/\varepsilon)$ that outputs an element of $G$ such that each $g\in G$ is output with probability in range $(1/|G|-\varepsilon,1/|G|+\varepsilon)$.
Theorem 2.4 ([8]).
Let $\mathcal{G}$ be a black-box group. For any $G\leq\mathcal{G}$, there exists a classical randomized algorithm running in time polynomial in $\log(|\mathcal{G}|)$ that decides if $G$ is solvable.
Watrous’ algorithms for solvable groups.
Watrous showed that for solvable groups, Group Order Verification and Group Non-Membership can be solved by polynomial-time quantum algorithms. Here are the precise statements we will use in our paper:
Theorem 2.5.
([43]) Let $\mathcal{G}$ be a black-box group. There exist quantum algorithms running in time $\mathrm{poly}(\log|\mathcal{G}|)$ that solve the following tasks with probability at least $1-1/\mathrm{poly}(|\mathcal{G}|)$:
  • given a solvable group $G\leq\mathcal{G}$, compute $|G|$;
  • given a solvable group $G\leq\mathcal{G}$ and an element $g\in\mathcal{G}$, decide if $g\in G$.
We immediately obtain the following corollary.
Corollary 2.1.
Let $\mathcal{G}$ be a black-box group. There exists a quantum algorithm running in time $\mathrm{poly}(\log|\mathcal{G}|)$ that given a group $G\leq\mathcal{G}$ and a solvable subgroup $H\leq G$, decide if $H$ is normal in $G$ with probability at least $1-1/\mathrm{poly}(|\mathcal{G}|)$.
Membership, normality, solvability and isomorphism certificates.
Babai and Szemerédi [12] showed that for any group $G$, any set of generators $S\subseteq G$ and any element $g\in G$, there exists a straight-line program over $S$ of length at most $(1+\log|G|)^{2}$ that generates $g$. This leads to the concept of membership certificate.
Definition 2.3.
Let $\mathcal{G}$ be a black-box group. For a group $G\leq\mathcal{G}$ given by a set of generators $S$ and an element $g\in G$, a certificate of membership of $g$ in $G$ is a straight-line program over $S$ of length at most $(1+\log|\mathcal{G}|)^{2}$ that generates $g$.
Since such a membership certificate can be verified in polynomial time using the group oracle, the existence of membership certificate established in [12] shows that Group Membership is in the class $\mathrm{NP}$. Next, we introduce the notion of normality certificate, which was also used in [12] for checking the normality of a subgroup.
Definition 2.4.
Let $\mathcal{G}$ be a black-box group. For any group $G\leq\mathcal{G}$ given as $G=\langle g_{1},\ldots,g_{t}\rangle$ and any subgroup $H\leq G$ given as $H=\langle h_{1},\ldots,h_{s}\rangle$, a normality certificate of $H$ in $G$ is a collection of membership certificates for the inclusions
$g_{j}h_{i}g_{j}^{-1}\in H,\textrm{ for each }i\in[s]\textrm{ and each }j\in[t].$ (4)
We now introduce the notion of solvability certificate [12], which is used for checking if a group is a solvable group of order dividing a given integer.
Definition 2.5.
Let $\mathcal{G}$ be a black-box group. For a group $G\leq\mathcal{G}$ and an integer $m$, a certificate that $G$ is a solvable group of order dividing $m$ is
  • a list of $s$ primes $(p_{1},\ldots,p_{s})$ such that $p_{1}\cdots p_{s}=m$ ;
  • a set of elements $g_{1},\ldots,g_{s}\in G$ along with certificates certifying membership in $G$ ;
  • for each $i\in[s]$, a normality certificate certifying that $\langle g_{1},\ldots,g_{i-1}\rangle\unlhd\langle g_{1},\ldots,g_{i}\rangle$ ;
  • for each $i\in[s]$, a certificate of the inclusion $g_{i}^{p_{i}}\in\langle g_{1},\ldots,g_{i-1}\rangle$ .
Babai and Szemerédi have shown (see [12, Theorem 11.4]) that the order of a black-box group is certifiable if all its composition factors are isomorphic to simple groups that have polylogarithmic-length presentations. We will need the following slightly different statement, which is proved by a similar technique.
Proposition 2.1.
Let $\mathcal{G}$ be a black-box group and $S$ be a simple group that has a polylogarithmic-length presentation. For any $G\leq\mathcal{G}$, the problem of testing if $G\cong S$ is in $\mathrm{NP}$.
Proof.
Let $\alpha_{1},\ldots\alpha_{s}$ be the generators of $S$, and $\mathcal{R}$ be the set of relations between those generators in the polylogarithmic-length presentation of $S$. Let $\{g_{1},\ldots,g_{t}\}$ denote the set of generators of $G$. If $G=\{e\}$ (which can be easily checked) we know that $G$ is not isomorphic to $S$. We thus assume below that $G\neq\{e\}$.
The prover sends elements $g^{\prime}_{1},\ldots,g^{\prime}_{s}$ of $G$, along with certificates of the equality $\langle g^{\prime}_{1},\ldots,g^{\prime}_{s}\rangle=\langle g_{1},\ldots,g_{t}\rangle$. The verifier checks if the certificates are correct (which guarantees that $\langle g^{\prime}_{1},\ldots,g^{\prime}_{s}\rangle=G$) and checks if each relation in $\mathcal{R}$ holds in $G$ when replacing $\alpha_{i}$ by $g^{\prime}_{i}$ for all $i\in[s]$ (which guarantees [^6] This is a folklore fact (see for instance [17, Section 6.3]). Here is a proof: the map $\varphi(\alpha_{i})=g^{\prime}_{i}$ can be extended into a surjective homomorphism $\varphi\colon S\to\langle g^{\prime}_{1},\ldots,g_{s}^{\prime}\rangle$. We thus have $\langle g^{\prime}_{1},\ldots,g_{s}^{\prime}\rangle\cong S/\ker{(\varphi)}$. that the group $\langle g^{\prime}_{1},\ldots,g_{s}^{\prime}\rangle$ is isomorphic to $S/N$ for some normal subgroup $N$ of $S$).
If $G$ is isomorphic to $S$, then there exists an isomorphism $\phi\colon S\to G$. The prover sends the elements $g^{\prime}_{1},\ldots,g^{\prime}_{s}$ of $G$ such that $\phi(\alpha_{i})=g^{\prime}_{i}$ for each $i\in[s]$. We have $\langle g^{\prime}_{1},\ldots,g^{\prime}_{s}\rangle=\langle g_{1},\ldots,g_{t}\rangle$, and thus the prover can also send correct certificates of this equality. Since $\phi$ is a homomorphism, each relation in $\mathcal{R}$ holds in $G$ when replacing $\alpha_{i}$ by $g^{\prime}_{i}$ for all $i\in[s]$. Thus all the tests succeed.
Conversely, if all the tests succeed we know that $G=\langle g^{\prime}_{1},\ldots,g^{\prime}_{s}\rangle$ is isomorphic to $S/N$ for some normal subgroup $N$ of $S$. Since $S$ is simple, its only normal subgroups are $\{e\}$ or $S$. Since the case $G=\{e\}$ is excluded, we conclude that $G$ is isomorphic to $S$. ∎
Using Proposition 2.1, we show the following result.
Proposition 2.2.
Let $\mathcal{G}$ be a black-box group and $\mathcal{S}$ be a multiset of simple groups that each has a polyloga-rithmic-length presentation and is given by its standard name. For any $G\leq\mathcal{G}$, the problem of testing if the multiset of composition factors of $G$ is $\mathcal{S}$ is in $\mathrm{NP}$.
Proof.
Merlin guesses a composition series of $G$ :
$\{e\}=H_{0}\unlhd H_{1}\unlhd\cdots\unlhd H_{s}=G\,.$
For each $i\in[s]$, Merlin sends to Arthur a set of generators for $H_{i}$, a normality certificate certifying that $H_{i-1}$ is normal in $H_{i}$ and the standard name $z_{i}$ of the simple group $H_{i}/H_{i-1}$. Arthur checks that the normality certificates are correct and also checks that $\{z_{1},\ldots,z_{s}\}$ is equal to the multiset of standard names corresponding to $\cal S$}. He then checks that $H_{i}/H_{i-1}\cong\mathrm{gr}(z_{i})$ for all $i\in[s]$ using the protocol from Proposition 2.1. ∎
3 The Babai-Beals Filtration and Nice Group Decompositions
In this section, we define our concept of nice decomposition of a group, inspired by the Babai-Beals filtration. We first describe the Babai-Beals filtration in Section 3.1. We then introduce the notion of nice decomposition and show several important properties in Section 3.2.
3.1 The Babai-Beals filtration
The following characteristic chain of subgroups introduced by Babai and Beals [8] has become a fundamental tool in the algorithmic theory of matrix and black-box groups: for any group $G$,
$\{e\}\leq\mathrm{Sol}(G)\leq\mathrm{Soc^{*}}(G)\leq\mathrm{Pker}(G)\leq G\,.$
We now define each of the terms in this chain.
  • $\mathrm{Sol}(G)$, the solvable radical of $G$, is the unique largest solvable normal subgroup of $G$.
  • ${\rm Soc}(G)$, the socle of $G$, is the subgroup generated by all minimal normal subgroups of $G$.
  • $\mathrm{Soc^{*}}(G)$ is the preimage of ${\rm Soc}(G/\mathrm{Sol}(G))$ in the natural projection $G\longrightarrow G/\mathrm{Sol}(G)$, i.e., $\mathrm{Soc^{*}}(G)/\mathrm{Sol}(G)={\rm Soc}(G/\mathrm{Sol}(G))$. The group $\mathrm{Soc^{*}}(G)/\mathrm{Sol}(G)={\rm Soc}(G/\mathrm{Sol}(G))$ is the direct product of simple groups $T_{1},\ldots,T_{k}$. The group $G$ acts on the set $\{T_{1},\ldots,T_{k}\}$ via conjugational action. Let $\phi:G\longrightarrow\mathrm{Sym}(k)$ denote the permutation representation of $G$ via conjugation action on the set $\{T_{1},\ldots,T_{k}\}$.
  • $\mathrm{Pker}(G)$, the permutation kernel of $G$, is the kernel of $\phi$, i.e., $\mathrm{Pker}(G)=\ker(\phi)$.
In this paper, we will not require detailed knowledge of all the algebraic structure of these terms. Instead, we will only need the following four properties ([8, Section 1.2]) :
  • (a)
    $\mathrm{Sol}(G)$ is solvable;
  • (b)
    $\mathrm{Soc^{*}}(G)/\mathrm{Sol}(G)$ is a direct product of simple groups;
  • (c)
    $\mathrm{Pker}(G)/\mathrm{Soc^{*}}(G)$ is solvable;
  • (d)
    $G/\mathrm{Pker}(G)\leq\mathrm{Sym}(k)$ and $k\leq\frac{\log|G|}{\log 60}$.
A set of generators of $\mathrm{Pker}(G)$ can be computed in Monte Carlo polynomial time [8, Theorem 1.1]. [^7] Theorem 1.1 in [8] assumes that a superset of the set of primes dividing $|G|$ is given as an additional input. These primes are used to compute the order of elements of $G$ efficiently. In the quantum setting, this is not needed since the order of an element can be computed in polynomial time [15, 40]. Moreover, in Monte Carlo polynomial time, we can set up a data structure which allows, for any $g\in G$, to compute $\phi(g)$ in deterministic polynomial time ([8, Corollary 5.2]). Using this data structure, membership in $\mathrm{Pker}(G)$ can checked in deterministic polynomial time (see the discussion after Theorem 1.1 in [8]). The following tasks can also be solved in deterministic polynomial time: [^8] As explained after Theorem 1.1 in [8], this can be done by using the extensive library of polynomial-time algorithms for permutation groups [11, 16, 35, 36]. More specifically, we can use the algorithms from [35, Section 6], [36, Section 3] or [16, Section 3.6].
  • compute a set of generators for $G/\mathrm{Pker}(G)\leq\mathrm{Sym}(k)$ as permutations in $\mathrm{Sym}(k)$;
  • compute the multiset of composition factors of $G/\mathrm{Pker}(G)$, where each composition factor is given by its standard name, and thus compute $|G/\mathrm{Pker}(G)|$.
3.2 Nice decompositions and their properties
We now introduce our concept of nice decomposition.
Definition 3.1.
Let $P$ be a group. A nice decomposition of $P$ consists of a solvable normal subgroup $H_{0}\unlhd P$ and $2s$ elements $\beta_{1},\ldots,\beta_{s},\gamma_{1},\ldots,\gamma_{s}\in P$ such that, when defining
$H_{i}$ $=\langle H_{0},\beta_{1},\ldots,\beta_{i},\gamma_{1},\ldots,\gamma_{i}\rangle$
for each $i\in[s]$, the following conditions are satisfied:
  • (C1)
    $\{e\}\leq H_{0}\unlhd H_{1}\unlhd\cdots\unlhd H_{s}\unlhd P$,
  • (C2)
    $P/H_{s}$ is solvable,
  • (C3)
    $\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}$ is simple, for all $i\in[s]$.
We first state a general property of nice decompositions that will be used in Section 5 to analyze the soundness of our protocol.
Proposition 3.1.
For any group $P$ and any nice decomposition of $P$ (with the same notations as in Definition 3.1), $|H_{i}/H_{i-1}|$ divides $|\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}|$ for any $i\in[s]$.
Proof.
Define $H^{\prime}=\langle H_{0},\beta_{i},\gamma_{i}\rangle\cap H_{i-1}$, which is a normal subgroup of $\langle H_{0},\beta_{i},\gamma_{i}\rangle$ since $H_{i-1}\unlhd H_{i}$. Observe that any element $g\in H_{i}$ can be written as
$g=h_{0}\,w_{1}(g)\cdots w_{i-1}(g)\,w_{i}(g),$ (5)
where $h_{0}\in H_{0}$ and $w_{\ell}(g)$ means one word over the alphabet $\{\beta_{\ell},\gamma_{\ell}\}$. This representation is not unique, but if $g$ also admits the representation
$g=h^{\prime}_{0}\,w^{\prime}_{1}(g)\cdots w^{\prime}_{i-1}(g)\,w^{\prime}_{i}(g),$
we have $w_{i}(g){w^{\prime}_{i}}^{-1}(g)\in\langle\beta_{i},\gamma_{i}\rangle\cap H_{i-1}\leq H^{\prime}$.
Define the map $\phi:H_{i}\longrightarrow\frac{\langle H_{0},\beta_{i},\gamma_{i}\rangle}{H^{\prime}}$ using Eq. (5) as
$\phi\big(h_{0}\,w_{1}(g)\cdots w_{i-1}(g)\,w_{i}(g)\big)=w_{i}(g)H^{\prime}$
(the above observation guarantees that the map is well-defined). This is a homomorphism: for any elements $g,g^{\prime}\in H_{i}$ written as
$g$ $=h_{0}\,w_{1}(g)\cdots w_{i-1}(g)\,w_{i}(g),$
$g^{\prime}$ $=h^{\prime}_{0}\,w^{\prime}_{1}(g)\cdots w^{\prime}_{i-1}(g)\,w^{\prime}_{i}(g),$
we obtain
$\phi(gg^{\prime})=w_{i}(g)w_{i}(g^{\prime})H^{\prime}=\phi(g)\phi(g^{\prime}).$
We have
$\ker(\phi)=\left\{h_{0}\,w_{1}(g)\cdots w_{i-1}(g)\,w_{i}(g)\>|\>w_{i}(g)\in H^{\prime}\right\}=H_{i-1}.$
Clearly, $\phi$ is surjective. By the first homomorphism theorem of groups, we conclude that
$\frac{H_{i}}{{H_{i-1}}}\cong\frac{\langle H_{0},\beta_{i},\gamma_{i}\rangle}{H^{\prime}}.$
Since $H_{0}\leq H^{\prime}$, we conclude that $|H_{i}/H_{i-1}|$ divides $|\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}|$. ∎
By using the Babai-Beals filtration we can prove the following theorem, which shows that $\mathrm{Pker}(G)$ has a nice decomposition with a very useful additional property (Property ($\star$)). This property will be crucial in Section 5 to establish the completeness of our protocol.
Theorem 3.1.
For any group $G$, the subgroup $\mathrm{Pker}(G)$ has a nice decomposition satisfying the condition
$H_{i}/H_{i-1}\cong\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}\,\textrm{, for all }i\in[s].$
Proof.
We use the Babai-Beals filtration. From the discussion in Section 3.1, there exists an isomorphism
$\phi\colon T_{1}\times\cdots\times T_{k}\to\mathrm{Soc^{*}}(G)/\mathrm{Sol}(G),$
for simple groups $T_{1},\ldots,T_{k}$. Since every simple group admits a generating set of size two (see, e.g., [22]), for each $i\in[k]$ we can write $T_{i}=\langle a_{i},b_{i}\rangle$. We take $s=k$ and $H_{0}=\mathrm{Sol}(G)$. For any $i\in[s]$, we take (arbitrary) elements $\beta_{i},\gamma_{i}\in\mathrm{Soc^{*}}(G)$ such that
$\phi((e,\ldots,e,a_{i},e\ldots,e))$ $=\beta_{i}H_{0}\,$,
$\phi((e,\ldots,e,b_{i},e\ldots,e))$ $=\gamma_{i}H_{0}\,$,
where in the above equations $a_{i}$ and $b_{i}$ are at the $i$-th position. For each $i\in[s]$, define the subgroup $H_{i}=\langle H_{0},\beta_{1},\ldots,\beta_{i},\gamma_{1},\ldots,\gamma_{i}\rangle$. Note that $H_{s}=\mathrm{Soc^{*}}(G)$ and $H_{i-1}\unlhd H_{i}$ for any $i\in[s]$. Additionally, $\mathrm{Pker}(G)/H_{s}=\mathrm{Pker}(G)/\mathrm{Soc^{*}}(G)$ is solvable, as explained in Section 3.1. Conditions (C1) and (C2) of Definition 3.1 thus hold for $P=\mathrm{Pker}(G)$. Since $H_{0}\unlhd G$, we also have $H_{0}\unlhd\mathrm{Pker}(G)$, as required for a nice decomposition.
Note that the following property holds:
$\langle H_{0},\beta_{i},\gamma_{i}\rangle\cap H_{i-1}=H_{0}\>\textrm{ for any }i\in[s].$ (6)
Observe that any element $g\in H_{i}$ can be written as
$g=h_{0}\,w_{1}(g)\cdots w_{i-1}(g)\,w_{i}(g),$
where $h_{0}\in H_{0}$ and $w_{\ell}(g)$ means one word over the alphabet $\{\beta_{\ell},\gamma_{\ell}\}$. This representation is not unique, but (6) shows that if $g$ also admits the representation
$g=h^{\prime}_{0}\,w^{\prime}_{1}(g)\cdots w^{\prime}_{i-1}(g)\,w^{\prime}_{i}(g),$
we have $w_{i}(g){w^{\prime}_{i}}^{-1}(g)\in\langle\beta_{i},\gamma_{i}\rangle\cap H_{i-1}\leq\langle H_{0},\beta_{i},\gamma_{i}\rangle\cap H_{i-1}=H_{0}$.
Define the map $\psi:H_{i}\longrightarrow\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}$ as
$\psi\big(h_{0}\,w_{1}(g)\cdots w_{i-1}(g)\,w_{i}(g)\big)=w_{i}(g)H_{0}\,.$
Clearly, $\psi$ is a surjective homomorphism and $H_{i-1}\leq\ker(\psi)$. Since $\ker(\psi)\neq H_{i}$, and $H_{i}/H_{i-1}\cong T_{i}$ is a simple group, $H_{i-1}=\ker(\psi)$. By the first homomorphism theorem of groups, we get $H_{i}/H_{i-1}\cong\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}$, i.e., Condition $(\star)$ holds. Combined with the fact that each $H_{i}/H_{i-1}$ is simple, this implies Condition (C3) of Definition 3.1. We conclude that $\mathrm{Pker}(G)$ has a nice decomposition satisfying Condition ($\star$). ∎
4 Testing Isomorphism to a Ree Group of Rank One
In this section, we study the following problem.
[…]
Input: $\ast$ a set of generators of a solvable subgroup $L\leq\mathcal{G}$ $\ast$ two elements $\beta,\gamma\in\mathcal{G}$ such that $L$ is normal in $\langle\beta,\gamma,L\rangle$ $\ast$ an integer $q$ of the form $3^{2a+1}$ for some $a>0$ Output: yes if $\langle\beta,\gamma,L\rangle/L$ is isomorphic to $\mathrm{R}(q)$; no otherwise
Here is the main result of this section.
Theorem 4.1.
[…]
We prove Theorem 4.1 in Section 4.2, after introducing in Section 4.1 tools to handle $\mathrm{R}(q)$ in its standard representation.
4.1 Constructive membership for the Ree group in its standard representation
Consider the standard representation $\mathrm{R}(q)=\langle\Gamma_{1},\Gamma_{2},\Gamma_{3}\rangle$ introduced in Section 2.2. We will need to implement constructive membership in this representation, i.e., given a matrix $M\in\mathrm{GL}(7,q)$ that belongs to $\langle\Gamma_{1},\Gamma_{2},\Gamma_{3}\rangle$, find a straight-line program over $\{\Gamma_{1},\Gamma_{2},\Gamma_{3}\}$ that reaches $M$. Since $\mathbb{F}_{q}$ has characteristic 3, we can use the following randomized algorithm for constructive membership in matrix groups of odd characteristic from [9]:
Theorem 4.2 (Theorem 2.3 in [9]).
There exists a randomized polynomial-time algorithm that solves the constructive membership problem in matrix groups of odd characteristic, given number-theoretic oracles.
The output of the randomized algorithm of Theorem 4.2 is not unique: the straight-line program output by the algorithm depends on the random bits used by the algorithm. Since in our applications we will need to specify a unique output, we consider the algorithm of Theorem 4.2 as a deterministic algorithm that receives as an auxiliary input a seed of random bits (which we denote by $\lambda$). The number-theoretic oracles referred to in Theorem 4.2 are oracles for integer factoring and discrete logarithm. Since these two tasks can be implemented in polynomial time using a quantum computer [40], this gives a polynomial-time quantum algorithm for constructive membership in $\mathrm{R}(q)$, which we denote by $\mathsf{ReeMemb}_{q}(\lambda)$. Here is the precise statement.
Corollary 4.1.
For any $q$, there exists a collection of polynomial-time quantum algorithms
$\Big\{\mathsf{ReeMemb}_{q}(\lambda)\>|\>\lambda\in\{0,1\}^{\mathrm{poly}(\log q)}\Big\}$
that receive as input a matrix $M\in\mathrm{GL}(7,q)$ and satisfy the following condition for any $M\in\mathrm{R}(q)$: For a fraction at least $1-10^{-7}$ of the $\lambda$’s, $\mathsf{ReeMemb}_{q}(\lambda)$ outputs “success” with probability at least $1-1/\mathrm{poly}(q)$. When $\mathsf{ReeMemb}_{q}(\lambda)$ outputs “success” it also outputs a straight-line program over $\{\Gamma_{1},\Gamma_{2},\Gamma_{3}\}$ reaching $M$. This straight-line program depends only on $M$ and $\lambda$ (i.e., it does not depend on the measurement outcomes of the quantum algorithm).
Let $\mathsf{Valid}(\lambda)$ denote the subset of $\mathrm{R}(q)$ containing all the $M\in\mathrm{R}(q)$ such that $\mathsf{ReeMemb}_{q}(\lambda)$ on input $M$ outputs “success” with probability at least $1-1/\mathrm{poly}(q)$, for the same polynomial as in Corollary 4.1. Let $\mathsf{Good}$ denote the $\lambda$’s such that
$\Pr_{M\in\mathrm{R}(q)}[M\in\mathsf{Valid}(\lambda)]\geq 1-10^{-5}\,.$
The following claim follows from Corollary 4.1 by a counting argument.
Claim 1.
When $\lambda$ is taken uniformly at random,
$\Pr[\lambda\in\mathsf{Good}]\geq 0.99\,.$
Proof.
Let $\delta=2^{\mathrm{poly}(\log q)}$ denote the total number of seeds $\lambda$. The number of pairs $(\lambda,M)$ such that $M\in\mathsf{Valid}(\lambda)$ is $(1-10^{-7})\delta|\mathrm{R}(q)|$. Note that for each $\lambda\notin\mathsf{Good}$ there are at most $(1-10^{-5})|\mathrm{R}(q)|$ pairs $(\lambda,M)$ with $M\in\mathsf{Valid}(\lambda)$, while for each $\lambda\in\mathsf{Good}$ there are (obviously) at most $|\mathrm{R}(q)|$ pairs $(\lambda,M)$ with $M\in\mathsf{Valid}(\lambda)$. We thus have
$|\mathsf{Good}||\mathrm{R}(q)|+(\delta-|\mathsf{Good}|)(1-10^{-5})|\mathrm{R}(q)|\geq(1-10^{-7})\delta|\mathrm{R}(q)|,$
which implies
$\frac{|\mathsf{Good}|}{\delta}\geq\frac{10^{-5}-10^{-7}}{10{{}^{-5}}}=0.99,$
as claimed. ∎
4.2 Proof of Theorem 4.1
We are now ready to give the proof of the main result of this section.
Proof of Theorem 4.1.
We write $K=\langle\beta,\gamma,L\rangle$. We use the standard representation $\mathrm{R}(q)=\langle\Gamma_{1},\Gamma_{2},\Gamma_{3}\rangle$ introduced in Section 2.2 and studied in Section 4.1.
Merlin’s witness.
Merlin sends three elements $g_{1},g_{2},g_{3}\in K$ along with certificates of the following memberships: $g_{i}\in K$ for all $i\in\{1,2,3\}$, $\beta\in\langle g_{1},g_{2},g_{3},L\rangle$ and $\gamma\in\langle g_{1},g_{2},g_{3},L\rangle$.
Definition of the isomorphism candidate.
Before explaining Arthur’s checking procedure, we introduce a map $f_{\lambda}\colon\mathrm{R}(q)\to K/L$ defined by the three elements $g_{1},g_{2},g_{3}$ sent by Merlin and a binary string $\lambda$ that will be later chosen by Arthur.
Given as input $x\in\mathsf{Valid}(\lambda)$, Algorithm $\mathsf{ReeMemb}_{q}(\lambda)$ described in Section 4.1 outputs with probability at least $1-1/\mathrm{poly}(q)$ a straight-line program $(w_{1},w_{2},\ldots,w_{s})$ over $\{\Gamma_{1},\Gamma_{2},\Gamma_{3}\}$ reaching $x$, where $s=\mathrm{polylog}(|\mathrm{R}(q)|)$. Remember that this means that $w_{s}=x$ and each $w_{i}$ is either
  • (i)
    a member of $\{\Gamma_{1},\Gamma_{2},\Gamma_{3}\}$, or
  • (ii)
    an element of the form $w_{j}^{-1}$ or $w_{j}w_{k}$ from some $j,k<i$.
From the output $(w_{1},w_{2},\ldots,w_{s})$ of $\mathsf{ReeMemb}_{q}(\lambda)$ on input $x$, we define a new straight-line program $(w^{\prime}_{1},w^{\prime}_{2},\ldots,w^{\prime}_{s})$ over $\{g_{1},g_{2},g_{3}\}$ by replacing $\Gamma_{1}$ by $g_{1}$, $\Gamma_{2}$ by $g_{2}$ and $\Gamma_{3}$ by $g_{3}$ in each Case (i). For instance, the straight-line program
$(w_{1}=\Gamma_{1},w_{2}=\Gamma_{2},w_{3}=w_{1}w_{1},w_{4}=w_{3}w_{2},w_{5}=w_{4}^{-1},w_{6}=\Gamma_{3},w_{7}=w_{5}w_{6})$
reaching the element $(\Gamma_{1}\Gamma_{1}\Gamma_{2})^{-1}\Gamma_{3}$ becomes the straight-line program
$(w^{\prime}_{1}=g_{1},w^{\prime}_{2}=g_{2},w^{\prime}_{3}=w^{\prime}_{1}w^{\prime}_{1},w^{\prime}_{4}=w^{\prime}_{3}w^{\prime}_{2},w^{\prime}_{5}={w^{\prime}}_{\!\!4}^{-1},w^{\prime}_{6}=g_{3},w^{\prime}_{7}=w^{\prime}_{5}w^{\prime}_{6})$
reaching the element $(g_{1}g_{1}g_{2})^{-1}g_{3}$. We denote by $g_{\lambda}(x)$ the element of $K$ reached by $(w^{\prime}_{1},w^{\prime}_{2},\ldots,w^{\prime}_{s})$.
Define the map $f_{\lambda}\colon\mathrm{R}(q)\to K/L$ as follows: for any $x\in\mathrm{R}(q)$,
$f_{\lambda}(x)=\left\{\begin{array}[]{cl}g_{\lambda}(x)L&\textrm{ if }x\in\mathsf{Valid}(\lambda)\,,\\ L&\textrm{ otherwise}\,.\end{array}\right$.
We state the following elementary, but crucial, property of this map.
Claim 2.
For any $g_{1},g_{2},g_{3}\in K$, if there exists a homomorphism $\varphi\colon\mathrm{R}(q)\to K/L$ such that $\varphi(\Gamma_{i})=g_{i}L$ for each $i\in\{1,2,3\}$, then $f_{\lambda}(x)=\varphi(x)$ for any $x\in\mathsf{Valid}(\lambda)$.
Arthur’s checking procedure.
Arthur’s main verification procedure is the procedure $\mathsf{IsoCheck}$ described below, which uses $\mathsf{Maj}(x)$ as a subprocedure. This verification procedure uses (at Section 4.2 of $\mathsf{IsoCheck}$ and Section 4.2 of $\mathsf{Maj}(x)$) the sampling algorithm from Theorem 2.3 (with $\varepsilon=1/\mathrm{poly}(|\mathcal{G}|)$) to sample a random element a constant number of times. As explained below, it also uses a constant number of times Watrous’ algorithm (Theorem 2.5) for membership in a solvable group. Since the sampling algorithm only performs approximate sampling and the second algorithm only succeeds with high probability, this may introduce some errors. These errors are nevertheless exponentially small, and thus have a negligible impact on the overall success probability. Additionally, when applying Algorithm $\mathsf{ReeMemb}_{q}(\lambda)$ on an input $M\in\mathsf{Valid}(\lambda)$, the error probability (which is exponential small) also has a negligible impact on the overall success probability. For simplicity, we will thus simply ignore all these failure probabilities in the discussion below.
$\mathsf{IsoCheck}$
// Checks if $K/L$ is isomorphic to $\mathrm{R}(q)$
1Choose $\lambda$ uniformly at random ;
2repeat $\mathrm{12\,times}$ :
3   Take two elements $r_{1}$ and $r_{2}$ uniformly at random from $\mathrm{R}(q)$;
4 if $g_{\lambda}(r_{1}r_{2})g_{\lambda}(r_{2})^{-1}g_{\lambda}(r_{1})^{-1}\notin L$ then output “no” ;
5
[…]
7
8if $\mathsf{Maj}(\Gamma_{i})=g_{i}L\,\,\mathrm{for\,all}\,\,i\in\{1,2,3\}$ then output “yes” ;
9 else output “no” ;
$\mathsf{Maj}(x)$
// computes $\phi(x)$ for the $\phi$ of Lemma 2.1
[…]
2for $i$ from $1$ to $s$ do
3   Take an element $r$ uniformly at random from $\mathrm{R}(q)$ ;
4 $h_{i}=g_{\lambda}(xr)g_{\lambda}(r^{-1})$ ;
5return the coset of $L$ that appears the most frequently among $h_{1}L,h_{2}L,\ldots,h_{s}L$ (breaking ties arbitrarily) ;
At Section 4.2 of $\mathsf{IsoCheck}$, to check if $K=L$ we only need to check if $\beta\in L$ and $\gamma\in L$, which can be done using Theorem 2.5. Checking if Merlin’s certificates are correct is straightforward. At Section 4.2, we compute $g_{\lambda}(r_{1})$, $g_{\lambda}(r_{2})$ and $g_{\lambda}(r_{1}r_{2})$ in polynomial time by decomposing $r_{1}$, $r_{2}$ and $r_{1}r_{2}$ using Algorithm $\mathsf{ReeMemb}_{q}(\lambda)$, compute $g_{\lambda}(r_{1}r_{2})g_{\lambda}(r_{2})^{-1}g_{\lambda}(r_{1})^{-1}$ using the black box for the group $\mathcal{G}$, and test membership in $L$. At Section 4.2, we use Theorem 2.5 to check if $\mathsf{Maj}(\Gamma_{i})=g_{i}L$.
At Section 4.2 of $\mathsf{Maj}$, we compute $g_{\lambda}(xr)$ and $g_{\lambda}(r^{-1})$ in polynomial time by using Algorithm $\mathsf{ReeMemb}_{q}(\lambda)$, and compute $g_{\lambda}(xr)g_{\lambda}(r^{-1})$ using the black box for the group $\mathcal{G}$. At Section 4.2 of $\mathsf{Maj}$, we use Theorem 2.5 to compare the cosets and select the one that appears the most frequently.
Completeness.
If $K/L\cong\mathrm{R}(q)$, then there exists an isomorphism $\varphi\colon\mathrm{R}(q)\to K/L$. We assume below that $\lambda\in\mathsf{Good}$, which happens with probability at least 0.99 (1).
Merlin sends $g_{1},g_{2},g_{3}\in K$ such that $g_{i}L=\varphi(\Gamma_{i})$ for each $i\in\{1,2,3\}$, as well as a correct certificate of the membership $g_{i}\in K$, for each $i\in\{1,2,3\}$. Since $\varphi$ is surjective, Merlin can also send correct certificates of the memberships $\beta\in\langle g_{1},g_{2},g_{3},L\rangle$ and $\gamma\in\langle g_{1},g_{2},g_{3},L\rangle$.
Since $\lambda\in\mathsf{Good}$, we know that $r_{1},r_{2}$ and $r_{1}r_{2}$ are all in $\mathsf{Valid}(\lambda)$ with probability at least 0.99997, in which case $f_{\lambda}(r_{1}r_{2})=g_{\lambda}(r_{1}r_{2})L$, $f_{\lambda}(r_{1})=g_{\lambda}(r_{1})L$ and $f_{\lambda}(r_{2})=g_{\lambda}(r_{2})L$ hold. 2 also implies that $f_{\lambda}(r_{1}r_{2})=\varphi(r_{1}r_{2})$, $f_{\lambda}(r_{1})=\varphi(r_{1})$ and $f_{\lambda}(r_{2})=\varphi(r_{2})$. Since $\varphi(r_{1}r_{2})=\varphi(r_{1})\varphi(r_{2})$, we obtain $g_{\lambda}(r_{1}r_{2})g_{\lambda}(r_{2})^{-1}g_{\lambda}(r_{1})^{-1}\in L$. We conclude that Procedure $\mathsf{IsoCheck}$ does not output “no” at Section 4.2 with probability at least $1-12\cdot 0.00003>0.99$.
From a similar argument, at Section 4.2 of $\mathsf{Maj}(x)$ we have $h_{i}L=f_{\lambda}(xr)f_{\lambda}(r^{-1})=\varphi(xr)\varphi(r^{-1})=\varphi(x)$ with probability at least $0.99998$. Among the 50 trials, we always get $h_{i}L=\varphi(x)$ with probability at least $0.999$, in which case we have $\mathsf{Maj}(x)=\varphi(x)$. In particular, we have $\mathsf{Maj}(\Gamma_{i})=\varphi(\Gamma_{i})=g_{i}L$ for all $i\in\{1,2,3\}$ with probability at least $(0.999)^{3}>0.99$.
The overall probability that $\mathsf{IsoCheck}$ outputs “yes” is thus at least $(0.99)^{3}>2/3$.
Soundness.
Now consider the case $K/L\not\cong\mathrm{R}(q)$. In the following, we assume that $\lambda\in\mathsf{Good}$, which happens with probability at least 0.99 (1). Assume for now that $\mathsf{Valid}(\lambda)=\mathrm{R}(q)$.
Consider first the case
$\Pr_{r_{1},r_{2}\in\mathrm{R}(q)}\>[f_{\lambda}(r_{1}r_{2})=f_{\lambda}(r_{1})f_{\lambda}(r_{2})]<9/10.$
In this case, $\mathsf{IsoCheck}$ outputs “no” at Section 4.2 at least once during the 12 iterations with probability at least $1-(9/10)^{12}>0.7$.
Now consider the case
$\Pr_{r_{1},r_{2}\in\mathrm{R}(q)}\>[f_{\lambda}(r_{1}r_{2})=f_{\lambda}(r_{1})f_{\lambda}(r_{2})]\geq 9/10.$
If $K=L$ or Merlin’s certificates are incorrect, $\mathsf{IsoCheck}$ outputs “no” at Section 4.2. We thus assume below that $K\neq L$ (i.e., $K/L\neq\{e\}$) and Merlin’s certificate are correct (i.e., $\langle g_{1},g_{2},g_{3},L\rangle=K$). Lemma 2.1 shows that there exists a homomorphism $\phi\colon\mathrm{R}(q)\to K/L$ such that for any $x\in\mathrm{R}(q)$,
$\Pr_{r\in\mathrm{R}(q)}\>[f_{\lambda}(xr)f_{\lambda}(r^{-1})=\phi(x)]$ $=\Pr_{r\in\mathrm{R}(q)}\>[f_{\lambda}(xr)f_{\lambda}(r^{-1})=\phi(xr)\phi(r^{-1})]$
$=1-\Pr_{r\in\mathrm{R}(q)}\>[f_{\lambda}(xr)f_{\lambda}(r^{-1})\neq\phi(xr)\phi(r^{-1})]$
$\geq 1-\Pr_{r\in\mathrm{R}(q)}\>[f_{\lambda}(xr)\neq\phi(xr)]-\Pr_{r\in\mathrm{R}(q)}\>[f_{\lambda}(r^{-1})\neq\phi(r^{-1})]$
$\geq 1-\frac{1}{10}-\frac{1}{10}$
$=\frac{8}{10}.$
Among 50 trials, the expected number of times we get $\phi(x)$ at Section 4.2 of $\mathsf{Maj}(x)$ is thus at least $(8/10)\cdot 50$. From Chernoff’s bound, this implies that $\phi(x)$ appears at least 26 times among the 50 times with probability at least
$1-\exp\left(-\frac{(7/20)^{2}\cdot(8/10)\cdot 50}{2}\right)>0.9,$
in which case we have $\mathsf{Maj}(x)=\phi(x)$. In particular,
$\mathsf{Maj}(\Gamma_{i})=\phi(\Gamma_{i})\textrm{ for all }i\in\{1,2,3\}$ (7)
holds with probability at least 0.7. If Eq. 7 holds, then there should be an index $i\in\{1,2,3\}$ such that $\mathsf{Maj}(\Gamma_{i})\neq g_{i}L$. Otherwise, $\phi$ would be a surjective homomorphism from $\mathrm{R}(q)$ to $K/L$, and thus an isomorphism since $\mathrm{R}(q)$ is a simple group and $K/L\neq\{e\}$, which contradicts the assumption $K/L\not\cong\mathrm{R}(q)$. The probability that $\mathsf{IsoCheck}$ outputs “no” at Section 4.2 is thus at least 0.7.
We actually have $\mathsf{Valid}(\lambda)\neq\mathrm{R}(q)$. The probability that the arguments of the $12\cdot 3+50\cdot 2=136$ calls to the function $g_{\lambda}$ performed by $\mathsf{IsoCheck}$ are all in the set $\mathsf{Valid}(\lambda)$ is nevertheless at least $1-132\cdot 10^{-5}>0.98$. The overall probability that $\mathsf{IsoCheck}$ outputs “no” is thus at least 0.70.010.02>2/30.7-0.01-0.02>2/3. ∎
5 Proof of Theorem 1.1
In this section, we prove Theorem 1.1, i.e., we show that checking that $|G|=m$ is in $\mathrm{QCMA}$, where $G=\langle g_{1},\ldots,g_{k}\rangle\leq\mathcal{G}$ and $m$ are the inputs of Group Order Verification (here $\mathcal{G}$ is a black-box group). We divide the proof into two parts: checking that $m$ divides $|G|$ (Proposition 5.1, which is the easy part) and checking that $|G|$ divides $m$ (Proposition 5.2, which is the hard part). Proposition 5.1 and Proposition 5.2 together immediately imply Theorem 1.1.
[…]
Adapting the classical strategy from [12, Section 9] to the quantum setting by replacing oracles for “independence testing” by efficient quantum algorithms dealing with solvable groups (Theorem 2.5), we obtain the following result.
Proposition 5.1.
There exists a $\mathrm{QCMA}$ protocol that checks if $m$ divides $|G|$.
Proof.
Merlin sends to Arthur the factorization of $m$ as a product of primes $m=p_{1}^{t_{1}}\cdots p_{r}^{t_{r}}$ and for each $i\in[r]$ a set of generators of a subgroup $G_{i}\leq G$, together with membership certificates certifying that $G_{i}$ is a subgroup of $G$. Arthur accepts if and only if the following three conditions hold:
  • (i)
    the factorization of $m$ is correct;
  • (ii)
    the membership certificates are correct;
  • (iii)
    $|G_{i}|=p_{i}^{t_{i}}$ for each $i\in[r]$.
Conditions (i) and (ii) can be checked in deterministic polynomial time. Since any group of order $p^{t}$ for some prime $p$ and some positive integer $t$ is solvable, Arthur can check Condition (iii) in quantum polynomial time by first checking solvability using Theorem 2.4 and then computing the order using Theorem 2.5.
The completeness and soundness of this protocol follow from 1. ∎
[…]
The hard part is to show that the order divides $m$. Combining all the techniques developed in this paper, we show the following result.
Proposition 5.2.
There exists a $\mathrm{QCMA}$ protocol that checks if $|G|$ divides $m$.
Proof.
We first describe the $\mathrm{QCMA}$ protocol and then analyze its correctness and soundness. Note that the completeness and soundness of the $\mathrm{QCMA}$ protocol of Theorem 4.1 can be amplified using standard techniques so that the completeness becomes $1-1/\mathrm{poly}(|\mathcal{G}|)$ and the soundness becomes $1/\mathrm{poly}(|\mathcal{G}|)$. In the following, we implicitly assume that the completeness and soundness have been amplified.
Merlin’s witness.
Merlin sends to Arthur a positive integer $m_{1}$ and elements $h_{1},\ldots,h_{n},k_{1},\ldots,k_{r}$, $\beta_{1},\ldots,\beta_{s}$, $\gamma_{1},\ldots,\gamma_{s}\in\mathcal{G}$, for some $n,r,s=O(\log|\mathcal{G}|)$. We write below
$H_{0}$ $=\langle h_{1},\ldots,h_{n}\rangle\,$,
$K$ $=\langle k_{1},\ldots,k_{r}\rangle\,$,
$H_{i}$ $=\langle H_{0},\beta_{1},\ldots,\beta_{i},\gamma_{1},\ldots,\gamma_{i}\rangle\>\textrm{ for each }i\in[s]\,.$
Merlin also sends to Arthur the following information:
  1. (i)
    membership certificates certifying that $h_{1},\ldots,h_{n},\beta_{1},\ldots,\beta_{s}$, $\gamma_{1},\ldots,\gamma_{s}$ are elements in $K$ ;
  2. (ii)
    normality certificates certifying that $K\unlhd G$, $H_{s}\unlhd K$, and $H_{i-1}\unlhd H_{i}$ for each $i\in[s]$ ;
  3. (iii)
    a solvability certificate certifying that $K/H_{s}$ is a solvable group of order divising $m_{1}$ ;
  4. (iv)
    a multiset $\mathcal{S}$ of simple groups with polylogarithmic-size presentation (each simple group is given by its standard name) and the certificate from Proposition 2.2 certifying that $\mathcal{S}$ is the set of composition factors of $G/K$ ;
  5. (v)
    for each $i\in[s]$, the standard name $z_{i}$ of a simple group ;
  6. (vi)
    for each $i\in[s]$ such that $z_{i}$ corresponds to the standard name of a finite simple group other than a Ree group of rank one, the certificate from Proposition 2.1 certifying that $\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}$ is isomorphic to $\mathrm{gr}(z_{i})$ ;
  7. (vii)
    for each $i\in[s]$ such that $z_{i}$ corresponds to the standard name of some Ree group of rank one $\mathrm{R}(q)$, the certificate from Theorem 4.1 certifying that $\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}$ is isomorphic to $\mathrm{R}(q)$ .
Terminology.
We use the following terminology below: we say that a certificate in (iv) or (vi) is correct if the checking procedure of Proposition 2.2 or Proposition 2.1, respectively, outputs “yes” on this certificate. We say that the certificate in (vii) is correct if the checking procedure of Theorem 4.1 outputs “yes” on this certificate with probability at least $1-1/\mathrm{poly}(|\mathcal{G}|)$. Note that for Proposition 2.1, Proposition 2.2 and Theorem 4.1, the existence of a correct certificate guarantees that the input is a yes-instance.
Arthur’s checking procedure.
Arthur first computes a set of generators of $\mathrm{Pker}(G)$, the multiset of composition factors of $G/\mathrm{Pker}(G)$ and the order $|G/\mathrm{Pker}(G)|$, which can be done in randomized polynomial time as discussed in Section 3.1. We write $m_{2}=|G/\mathrm{Pker}(G)|$. Arthur then checks that
  1. (1)
    $H_{0}$ is a solvable normal subgroup of $\mathrm{Pker}(G)$ ;
  2. (2)
    $k_{1},\ldots,k_{r}\in\mathrm{Pker}(G)$ ;
  3. (3)
    $\mathcal{S}$ is the multiset of composition factors of $G/\mathrm{Pker}(G)$ ;
  4. (4)
    each membership certificate in (i) is correct ;
  5. (5)
    each normality certificate in (ii) is correct, i.e., it certifies all the inclusions of Eq. 4 ;
  6. (6)
    the solvability certificate in (iii) is correct, i.e., it certifies all the conditions of Definition 2.5 ;
  7. (7)
    the certificate in (iv) is a correct certificate for Proposition 2.2;
  8. (8)
    the certificates in (vi) are correct certificates for Proposition 2.1;
  9. (9)
    the certificates in (vii) are correct certificates for Theorem 4.1;
  10. (10)
    the product $|H_{0}|\cdot|\mathrm{gr}(z_{1})|\cdots|\mathrm{gr}(z_{s})|\cdot m_{1}\cdot m_{2}$ divides $m$ .
[…]
Remark 1.
We cannot ask directly Merlin to prove that $K=\mathrm{Pker}(G)$ since Merlin does not know the elements of the generating set of $\mathrm{Pker}(G)$ computed by Arthur (generators of $\mathrm{Pker}(G)$ can only be computed in randomized polynomial time, not in deterministic polynomial time). Instead, we ask Merlin to send $\mathcal{S}$ and check that $K=\mathrm{Pker}(G)$ using Tests (2), (3) and (7).
Remark 2.
Even when $K/H_{s}$ is solvable, we cannot use Watrous’ quantum algorithm (Theorem 2.5) to compute its order since the group $K/H_{s}$ does not have unique encoding. This is why we ask Merlin to certify that the order of $K/H_{s}$ divides $m_{1}$ using the solvability certificate (iii), which can be checked classically even without unique encoding.
Completeness.
Assume that $|G|$ divides $m$. From Theorem 3.1, there exist a normal subgroup $H_{0}\unlhd\mathrm{Pker}(G)$ and $2s$ elements $\beta_{1},\ldots,\beta_{s},\gamma_{1},\ldots,\gamma_{s}\in\mathrm{Pker}(G)$ such that, when defining
$H_{i}$ $=\langle H_{0},\beta_{1},\ldots,\beta_{i},\gamma_{1},\ldots,\gamma_{i}\rangle$
for each $i\in[s]$, Conditions (C1), (C2), (C3) of Definition 3.1 and Condition ($\star$) of Theorem 3.1 are satisfied. We have
$|G|=|H_{0}|\cdot|H_{1}/H_{0}|\cdots|H_{s}/H_{s-1}|\cdot|\mathrm{Pker}(G)/H_{s}|\cdot|G/\mathrm{Pker}(G)|.$
Merlin sends generators of this subgroup $H_{0}$, generators of $K=\mathrm{Pker}(G)$, these $2s$ elements $\beta_{1},\ldots,\beta_{s},\gamma_{1},\ldots,\gamma_{s}$, and the integer $m_{1}=|\mathrm{Pker}(G)/H_{s}|$. Merlin sends the multiset $\mathcal{S}$ of composition factors of $G/\mathrm{Pker}(G)$. Each $z_{i}$ sent by Merlin is the standard name of the simple group $\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}$.
The existence of correct normality certificates for (ii) follows from the normality of $H_{i}$, $H_{s}$ and $\mathrm{Pker}(G)$. The existence of a correct solvability certificate for (iii) follows from Condition (C2) of Definition 3.1. The existence of a correct certificate for (iv) follows from Proposition 2.2 combined with Theorem 2.2 and Property (d) of the Babai-Beals filtration described in Section 3.1, which guarantee that the composition factors of $G/\mathrm{Pker}(G)$ (i.e., the simple groups in $\mathcal{S}$) have a short presentation. The existence of correct certificates of (vi) and (vii) follow from Proposition 2.1 and Theorem 4.1, respectively.
[…]
$|G|=|H_{0}|\cdot|\mathrm{gr}(z_{1})|\cdots|\mathrm{gr}(z_{s})|\cdot m_{1}\cdot m_{2}.$
Since $|G|$ divides $m$, this quantity divides $m$, and thus checking Item (10) also succeeds with high probability.
Soundness.
[…]
Item (1) guarantees that $H_{0}$ is a solvable normal subgroup of $\mathrm{Pker}(G)$. Item (2) guarantees that $K$ is a subgroup of $\mathrm{Pker}(G)$. Items (3) and (7) guarantee that the multiset of composition factors of $G/K$ matches the multiset of composition factors of $G/\mathrm{Pker}(G)$, which implies $|K|=|\mathrm{Pker}(G)|$. We thus have $K=\mathrm{Pker}(G)$.
Item (4) and (5) further guarantee that
$\{e\}\unlhd H_{0}\unlhd H_{1}\unlhd\cdots\unlhd H_{s}\unlhd\mathrm{Pker}(G)\unlhd G.$
Item (6) guarantees that $|\mathrm{Pker}(G)/H_{s}|$ divides $m_{1}$. Item (8) and (9) guarantee that $\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}\cong\mathrm{gr}(z_{i})$ for each $i\in[s]$. Proposition 3.1 then implies that $|H_{i}/H_{i-1}|$ also divides $|\mathrm{gr}(z_{i})|$, for each $i\in[s]$. We conclude that $|G|$ must divide the quantity
$|H_{0}|\cdot|\mathrm{gr}(z_{1})|\cdots|\mathrm{gr}(z_{s})|\cdot m_{1}\cdot m_{2}.$
Since $|G|$ does not divide $m$, this implies that Item (10) must fail whenever the computation of $|H_{0}|$ is correct, which happens with high probability. ∎
6 Proofs of the Other Results
In this section, we discuss how to derive the other results of Table 1.
We first give the formal definition of the eight problems introduced in Section 1.2.
Group Isomorphism
Instance: Elements $g_{1},\,\ldots,\,g_{k}$ in some group $\mathcal{G}$, elements $h_{1},\,\ldots,\,h_{\ell}$ in some group $\mathcal{H}$.
Question: Are $\langle g_{1},\,\ldots,\,g_{k}\rangle$ and $\langle h_{1},\,\ldots,\,h_{\ell}\rangle$ isomorphic?
Homomorphism
Instance: Elements $g_{1},\,\ldots,\,g_{k}$ in some group $\mathcal{G}$, elements $h_{1},\,\ldots,\,h_{k}$ in some group $\mathcal{H}$.
Question: Is there a homomorphism $\phi:\langle g_{1},\,\ldots,\,g_{k}\rangle\to\langle h_{1},\,\ldots,\,h_{k}\rangle$ such that $\phi(g_{i})=h_{i}$
                                  for each $i\in[k]$?
Minimal Normal Subgroup
Instance: Elements $g_{1},\,\ldots,\,g_{k}$ and $h_{1},\,\ldots,\,h_{\ell}$ in some group $\mathcal{G}$.
Question: Is $\langle h_{1},\ldots,h_{\ell}\rangle$ a minimal normal subgroup of $\langle g_{1},\ldots,g_{k}\rangle$?
Proper Subgroup
Instance: Elements $g_{1},\,\ldots,\,g_{k}$ and $h_{1},\,\ldots,\,h_{\ell}$ in some group $\mathcal{G}$.
Question: Is $\langle h_{1},\>\ldots,\>h_{\ell}\rangle\,$ a proper subgroup of $\langle g_{1},\ldots,g_{k}\rangle$?
Simple Group
Instance: Elements $g_{1},\ldots,g_{k}$ in some group $\mathcal{G}$.
Question: Is $\langle g_{1},\ldots,g_{k}\rangle$ a simple group?
Intersection
Instance: Elements $g_{1},\>\ldots,\>g_{k}$, $h_{1},\>\ldots,\>h_{\ell}$, and $a_{1},\ldots,a_{t}$ in some group $\mathcal{G}$.
Question: Is $\langle a_{1},\ldots,a_{t}\rangle$ equal to the intersection of $\langle g_{1},\ldots,g_{k}\rangle$ and $\langle h_{1},\ldots,h_{\ell}\rangle$?
Centralizer
Instance: Elements $g_{1},\ldots,g_{k}$, $h_{1},\ldots,h_{\ell}$ and $a$ in some group $\mathcal{G}$.
Question: Is $\langle h_{1},\ldots,h_{\ell}\rangle$ equal to the centralizer of $a$ in $\langle g_{1},\ldots,g_{k}\rangle$?
Maximal Normal Subgroup
Instance: Elements $g_{1},\,\ldots,\,g_{k}$ and $h_{1},\,\ldots,\,h_{\ell}$ in some group $\mathcal{G}$.
Question: Is $\langle h_{1},\ldots,h_{\ell}\rangle$ a maximal normal subgroup of $\langle g_{1},\ldots,g_{k}\rangle$?
We give below the proofs of Corollaries 1.3, 1.4 and 1.5.
Corollary 1.3 (repeated).
Group Isomorphism is in the complexity class $\mathrm{QCMA}$.
Proof.
Group Isomorphism can be reduced to the Group Order Verification as follows (see, [7, Proposition 4.9]): Merlin guesses $k$ elements $h^{\prime}_{1},\,\ldots,\,h^{\prime}_{k}$ from $\langle h_{1},\,\ldots,\,h_{\ell}\rangle$ such that $h^{\prime}_{i}=\phi(g_{i})$ for each $i\in[k]$, for some isomorphism $\phi\colon\langle g_{1},\,\ldots,\,g_{k}\rangle\to\langle h_{1},\,\ldots,\,h_{\ell}\rangle$. Merlin also guesses $m=|\langle g_{1},\,\ldots,\,g_{k}\rangle|$, as well as membership certificates certifying that $\langle h^{\prime}_{1},\,\ldots,\,h^{\prime}_{k}\rangle=\langle h_{1},\,\ldots,\,h_{\ell}\rangle$.
Consider the subgroup $K=\langle(g_{1},h^{\prime}_{1}),\cdots,(g_{k},h^{\prime}_{k})\rangle$ of the group $\langle g_{1},\,\ldots,\,g_{k}\rangle\times\langle h_{1},\,\ldots,\,h_{\ell}\rangle$. Arthur checks that the membership certificates are correct and checks that $m=|\langle g_{1},\,\ldots,\,g_{k}\rangle|=|\langle h_{1},\,\ldots,\,h_{\ell}\rangle|=|K|$ using the $\mathrm{QCMA}$ protocol for Group Order Verification. ∎
Corollary 1.4 (repeated).
Homomorphism, Minimal Normal Subgroup, Proper Subgroup and Simple Group are in the complexity class $\mathrm{QCMA}\cap\mathrm{coQCMA}$.
Proof.
Homomorphism and Minimal Normal Subgroup can be reduced to Group Order in polynomial time (see, [7, Corollary 12.1] and its proof). The claim thus follows from Corollary 1.1.
To show that Proper Subgroup is in $\mathrm{QCMA}$, we follow the strategy of [42, Section 5]. Merlin guesses an element $a\in\langle g_{1},\,\ldots,\,g_{k}\rangle$ such that $a\not\in\langle h_{1},\,\ldots,\,h_{\ell}\rangle$ and membership certificates certifying that $a\in\langle g_{1},\,\ldots,\,g_{k}\rangle$ and $h_{i}\in\langle g_{1},\,\ldots,\,g_{k}\rangle$ for each $i\in[\ell]$. Arthur checks that the membership certificates are correct and checks that $a\not\in\langle h_{1},\,\ldots,\,h_{\ell}\rangle$ using the $\mathrm{QCMA}$ protocol for Group Non-Membership.
To show that Proper Subgroup is in $\mathrm{coQCMA}$, we observe that $\langle h_{1},\,\ldots,\,h_{\ell}\rangle$ is not a proper subgroup of $\langle g_{1},\,\ldots,\,g_{k}\rangle$ if and only if either $\langle h_{1},\,\ldots,\,h_{\ell}\rangle=\langle g_{1},\,\ldots,\,g_{k}\rangle$ or there exists an element $h\in\langle h_{1},\,\ldots,\,h_{\ell}\rangle$ such that $h\notin\langle g_{1},\,\ldots,\,g_{k}\rangle$. Merlin guesses which case holds. In the first case, he also guesses membership certificates certifying that $\langle h_{1},\,\ldots,\,h_{\ell}\rangle=\langle g_{1},\,\ldots,\,g_{k}\rangle$. In the second case, he also guesses an element $h\in\langle h_{1},\,\ldots,\,h_{\ell}\rangle$ such that $h\notin\langle g_{1},\,\ldots,\,g_{k}\rangle$. Arthur checks that the membership certificates are correct and checks that $h\notin\langle g_{1},\,\ldots,\,g_{k}\rangle$ using the $\mathrm{QCMA}$ protocol for Group Non-Membership.
Watrous [42, Section 5] showed that Simple Group is in $\mathrm{coQMA}$ by using a quantum proof for Group Non-Membership. From Corollary 1.2, we can conclude that Simple Group is in $\mathrm{coQCMA}$. Theorem 2.1, Proposition 2.1 and Theorem 4.1 together imply that Simple Group is in $\mathrm{QCMA}$. ∎
Corollary 1.5 (repeated).
Intersection, Centralizer and Maximal Normal Subgroup are in the complexity class $\mathrm{coQCMA}$.
Proof.
Watrous [42] showed that Intersection, Centralizer and Maximal Normal Subgroup are in $\mathrm{coQMA}$ by using a quantum proof for Group Non-Membership along with classical proofs for various other properties (see, [42, Section 5]). By Corollary 1.2, Group Non-Membership is in $\mathrm{QCMA}$. This implies that Intersection, Centralizer and Maximal Normal Subgroup are in $\mathrm{coQCMA}$. ∎
Acknowledgments
The authors are grateful to Scott Aaronson, Michael Levet and James Wilson for helpful correspondence, and to Hirotada Kobayashi for many fruitful discussions. The authors are supported by JSPS KAKENHI 22H00522, 24H00071, 24K22293, MEXT Q-LEAP JPMXS0120319794, JST CREST JPMJCR24I4 and JST ASPIRE JPMJAP2302.